OpenStudy (anonymous):

Write a system of two equations with the variables m and f in each to represent the following problem: To skate at Roller Heaven, each person must pay a monthly membership fee (m) and a fee for each session (f). In January, Art attended 5 sessions and paid $25 in all. In July, he attended 10 sessions, and he paid $40 in all.

OpenStudy (mertsj):

25=m+5f 40=m+10f